# Zihan Tan    Suyuan Huang    Guancheng Wan    Wenke Huang    He Li    Mang Ye

### v2

@@ 4.2 Node Label Information Reinforcement
First of all, we introduce the Structure-Aware Label Centrality (SALC) metric, denoted as $\Lambda^{\text{SALC}}_{u}$. It is defined as the combination of the Label Influence Centrality $\Lambda^{l}_{u}$ and the Structural Prominence Score $\Lambda^{s}_{u}$:
@@
where $\tilde{P}_{u,v}$ is the $(u,v)$-th element of the standard Personalized PageRank (PPR) matrix $\mathbf{\tilde{P}}$, and $\tau_{v}$ represents the prior importance score of node $v$, typically initialized to 1 for all nodes. The Structural Prominence Score captures the maximum influence exerted by any node on node $u$, weighted by its importance.
As clarified, the PPR matrix used for computing Label Influence Centrality is denoted as $\mathbf{\tilde{P}}^{(L)}$, and defined as:
$\mathbf{\tilde{P}}^{(L)}=\alpha\left(\mathbf{I}-(1-\alpha)\mathbf{D}^{-1}\mathbf{A^{\prime}}\right)^{-1}.$
Here, $\mathcal{V}_{L}$ denotes the set of labeled nodes, and $\tilde{P}^{(L)}_{v,u}$ represents the influence of node $v$ on node $u$ as captured by $\mathbf{\tilde{P}}^{(L)}$. To accurately capture the influence of labeled nodes, the inclusion of self-loops in $\mathbf{A^{\prime}}$ ensures that each labeled node’s own label contributes to its $\Lambda^{l}_{u}$. Compared to the intuitive approach of directly selecting labeled nodes, the SALC metric $\Lambda^{\text{SALC}}_{u}$ considers both structural representativeness of the nodes and diffusion of label signals, thus avoiding biases caused by isolated labeled nodes. It is also capable of selecting unlabeled nodes that still possess rich label signals and structural advantages. This improves knowledge quality, enriches the repository, and mitigates the LSD problem. After computing the SALC scores for all nodes, we rank the nodes based on their $\Lambda^{\text{SALC}}_{u}$ values and select the top $K$ nodes, where the default value of $K$ is $1/3$ of the total number of nodes. Subsequently, for each class $c$, the local prototype $\mathbf{H}_{c}^{i}$ at each client is computed as the mean feature vector of the selected nodes belonging to class $c$:
@@
where $\mathcal{V}_{c}^{i}$ represents the set of nodes categorized as class $c$ on client $i$, and $\mathbf{h}_{u}^{i}$ is the feature vector of node $u$ on client $i$. Once the local prototypes are computed, clients upload their prototypes to the server along with the node count. For each class $c$, the server aggregates the prototypes from $\alpha$ percent of the clients by weighting each local prototype according to its sample size. Four global anchor prototypes are constructed for each class. Each global prototype $\mathbf{H}_{c}^{g,k}$ is computed as:
@@
where $\mathbf{H}_{c}^{g,k}$ represents the $k$-th global prototype for class $c$, $\mathcal{N}_{c}^{k}$ is the set of clients randomly selected to contribute to the $k$-th global prototype for class $c$, and $C$ is the number of classes. The global repository $\mathcal{H}$ contains all the global prototypes and will be broadcast back to clients. Once the global knowledge repository is constructed, it is distributed to the local clients along with the model parameters. The global features $\mathbf{h}_{u}^{g}$ used in the following loss formulation refer to the frozen inference features extracted locally using the distributed global model. To guide local training, we introduce a federated knowledge distillation loss function designed to align local feature representations with the global prototypes:
@@
where $\text{KL}(\cdot,\cdot)$ represents the Kullback-Leibler divergence, and $\sigma(\cdot)$ is the softmax function applied to similarity scores computed by the similarity function $\varphi(\mathbf{h},\mathcal{H})$, which returns a vector of cosine similarities between the feature $\mathbf{h}$ and all global prototypes.
@@ 4.3 Frequency-aware Graph Modeling Alignment
Reconstruct. As defined in Sec. 3, the feature vector matrix of graph $\mathcal{G}$ is denoted as $\mathbf{H}$. In order to emphasize the local structure and more accurately capture meaningful similarities between nodes, we use the original feature matrix $\mathbf{H}$. For each node $u$, we identify its $k_{\text{sim}}$ most similar neighbors based on the cosine similarity of their feature vectors $\mathbf{h}_{u}$. We then construct a sparse self-similarity matrix $\mathbf{S^{\prime}}$ as:
@@
To characterize the topological structure between nodes, we calculate the graph Laplacian matrix $\mathbf{L^{\prime}}$ based on this sparse similarity matrix $\mathbf{S^{\prime}}$ as:
@@
where $\mathbf{D^{\prime}}$ is the diagonal degree matrix of $\mathbf{S^{\prime}}$.
Projection. We perform eigendecomposition on the Laplacians $\mathbf{L^{\prime}}^{i}$ and $\mathbf{L^{\prime}}^{g}$. For $\mathbf{L^{\prime}}^{i}$, let $\{\mathbf{u}^{\text{low},i}_{m}\}_{m=1}^{k_{\text{eig}}}$ be the eigenvectors corresponding to the smallest eigenvalues (low-frequency), and $\{\mathbf{u}^{\text{high},i}_{m}\}_{m=1}^{k_{\text{eig}}}$ for the largest eigenvalues (high-frequency). Similarly, for $\mathbf{L^{\prime}}^{g}$, we obtain $\{\mathbf{u}^{\text{low},g}_{m}\}_{m=1}^{k_{\text{eig}}}$ and $\{\mathbf{u}^{\text{high},g}_{m}\}_{m=1}^{k_{\text{eig}}}$. The feature matrix $\mathbf{H}$ is then projected onto each of these eigenvectors. For instance:
@@
Applying these projections for each $m\in\{1,\dots,k_{\text{eig}}\}$, we obtain several sets of projected feature matrices used in the loss computation. Specifically, local features $\mathbf{H}^{i}$ are projected onto low/high-frequency eigenvectors of the local graph, yielding $\mathbf{Z}^{i,\text{low}}_{m}$ and $\mathbf{Z}^{i,\text{high}}_{m}$, respectively. Similarly, frozen global inference features $\mathbf{H}^{g}$ are projected onto corresponding eigenvectors from $\mathbf{L^{\prime}}^{g}$, yielding $\mathbf{Z}^{g,\text{low}}_{m}$ and $\mathbf{Z}^{g,\text{high}}_{m}$. Loss $\mathcal{L}_{\text{{FGMA}}}$ is then defined as the sum of MSE over all eigenvector-projected pairs:
@@
This loss addresses spectral heterogeneity by aligning client and global signal characteristics in both spectral domains.
Overall Objective. Combining strategies Node Label Information Reinforcementand Frequency-aware Graph Modeling Alignment, our framework $S^{2}$reinforces label knowledge during local modeling and mitigates spectral heterogeneity. The final objective is:
@@
where $\mathcal{L}_{\text{CE}}$ denotes the standard cross-entropy loss for node classification, while $\lambda_{1}$ and $\lambda_{2}$ are balancing coefficients for the proposed methods.
@@ 5.1 Experimental Setup
Table 1: Comparison with the state-of-the-art methods on homophilic and heterophilic graph datasets. We report node classification accuracies and each value difference from FedAvg with the best result in bold.
@@
FedAvg [ASTAT17] $81.9\pm 0.7$ $74.3\pm 0.4$ $87.3\pm 0.3$ $72.8\pm 2.2$ $77.6\pm 2.7$ $79.6\pm 0.1$
FedProx [arXiv18] $82.1\pm 0.5$0.2 $74.4\pm 0.3$0.1 $87.9\pm 0.4$0.6 $73.5\pm 3.7$0.7 $77.3\pm 3.4$0.3 $79.7\pm 0.1$0.1
FedNova [NeurIPS20] $81.6\pm 1.2$0.3 $74.4\pm 0.4$0.1 $88.2\pm 0.5$0.9 $73.0\pm 4.4$0.2 $77.4\pm 4.2$0.2 $79.9\pm 0.4$0.3
FedFa [ICLR23] $82.7\pm 0.5$0.8 $74.9\pm 0.6$0.6 $87.8\pm 0.5$0.5 $73.9\pm 3.6$1.1 $78.1\pm 4.6$0.5 $80.1\pm 0.3$0.5
FedSage+ [NeurIPS19] $82.3\pm 0.7$0.4 $75.2\pm 0.3$0.9 $88.2\pm 0.7$0.9 $73.7\pm 4.0$0.9 $\mathbf{79.0\pm 3.3}$1.4 $79.9\pm 0.2$0.3
FedStar [AAAI23] $82.6\pm 0.5$0.7 $74.5\pm 0.3$0.2 $88.1\pm 0.6$0.8 $74.3\pm 2.7$1.5 $78.3\pm 4.7$0.7 $79.8\pm 0.1$0.2
FedPub [ICML23] $82.3\pm 0.8$0.4 $74.8\pm 0.7$0.5 $88.0\pm 0.4$0.7 $73.4\pm 3.5$0.6 $77.8\pm 3.1$0.2 $79.9\pm 0.2$0.3
FGSSL [IJCAI23] $82.6\pm 0.4$0.7 $74.9\pm 0.2$0.6 $87.6\pm 0.7$0.3 $73.6\pm 4.6$0.8 $77.8\pm 3.8$0.2 $79.9\pm 0.2$0.3
FedGTA [VLDB24] $82.4\pm 0.8$0.5 $75.1\pm 0.5$0.8 $87.7\pm 0.9$0.4 $72.6\pm 4.2$0.2 $77.8\pm 4.1$0.2 $80.2\pm 0.3$0.6
FGGP [AAAI24] $82.5\pm 0.4$0.6 $74.7\pm 0.5$0.4 $87.5\pm 0.4$0.2 $73.6\pm 2.8$0.8 $78.2\pm 3.4$0.6 $80.4\pm 0.3$0.8
$S^{2}$ (ours) $\mathbf{83.4\pm 0.2}$1.5 $\mathbf{76.0\pm 0.3}$1.7 $\mathbf{88.6\pm 0.2}$1.3 $\mathbf{74.8\pm 2.3}$2.0 $\mathbf{79.0\pm 1.0}$1.4 $\mathbf{80.5\pm 0.1}$0.9
5.1.1 Datasets
  • Cora (McCallum et al., 2000) dataset consists of 2708 scientific publications classified into one of seven classes. There are 5429 edges in the network of citations. 1433 distinct words make up the dictionary.
  • Citeseer (Giles et al., 1998) dataset consists of 3312 scientific publications classified into one of six classes and 4732 edges. The dictionary contains 3703 unique words.
  • Pubmed (Sen et al., 2008) dataset consists of 19717 scientific papers on diabetes that have been categorized into one of three categories in the PubMed database. The citation network has 44338 edges. A word vector from a dictionary with 500 unique terms that is TF/IDF weighted is used to describe each publication in the dataset.
  • Texas and Wisconsin datasets are subsets of the WebKB dataset (Craven et al., 1998). The WebKB dataset was introduced in 1998, comprising web pages from the computer science departments of various universities, including the University of Texas and the University of Wisconsin. The dataset is commonly used for tasks such as webpage classification and link prediction, serving as a benchmark for evaluating machine learning models in graph-based learning scenarios.
  • Minesweeper (Baranovskiy et al., 2023) dataset is a synthetic graph dataset inspired by the Minesweeper game. In this dataset, the graph is structured as a regular 100x100 grid, where each node represents a cell connected to its neighboring nodes, except for edge nodes, which have fewer neighbors. The primary task is to predict which nodes contain mines. This dataset is commonly used to evaluate the performance of GNNs under heterophily.
5.1.2 Baselines
@@
We compare $S^{2}$ with several state-of-the-art approaches, including traditional federated learning methods such as FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), and FedNova (Wang et al., 2020); personalized federated learning algorithms like FedFa (Zhou & Konukoglu, 2023); federated graph learning approaches including FGSSL (Huang et al., 2023a) and FGGP (Wan et al., 2024a); as well as personalized federated graph learning methods such as FedSage+ (Zhang et al., 2021b), FedStar (Tan et al., 2023), FedPub (Baek et al., 2023), and FedGTA (Li et al., 2024). This comprehensive set of baseline methods spans various federated and graph learning paradigms, allowing us to evaluate the performance and advantages of our proposed $S^{2}$ across diverse scenarios.
5.1.3 Implement Details
@@
Following prevalent methodologies in FGL research, we employ the Louvain community detection algorithm to partition the graph into subgraphs assigned to different clients. For each dataset, we divide the nodes into training, validation, and testing sets with ratios of 60%, 20% and 20% respectively. Additionally, we simulate various collaborative scenarios by configuring the number of clients to 10 for Cora, Citeseer, Pubmed, and Minesweeper datasets, and 3 for Texas and Wisconsin datasets. The primary evaluation metric is the node classification accuracy on the clients’ test sets. We conduct each experiment five times and report the average accuracy from the last five communication epochs as the final performance. We conduct experiments with the ACM-GCN (Luan et al., 2022), which achieves a strong ability on both homophilic and heterophilic graph datasets.

### v3

@@ 4.2 Node Label Information Reinforcement
First of all, we introduce the Structure-Aware Label Centrality (SALC) metric, denoted as $\Lambda^{\text{SALC}}_{u}$ for node $u$. It is defined as the combination of the label influence centrality $\Lambda^{l}_{u}$ and the structural prominence score $\Lambda^{s}_{u}$:
@@
where $\tilde{P}_{u,v}$ is the $(u,v)$-th element of the standard PPR matrix $\mathbf{\tilde{P}}$, and $\tau_{v}$ represents the prior importance score of node $v$, typically initialized to 1 for all nodes. The structural prominence score captures the maximum influence exerted by any node on node $u$, weighted by its importance. As clarified, the PPR matrix used for computing label influence centrality is denoted as $\mathbf{\tilde{P}}^{(L)}$, and defined as:
$\mathbf{\tilde{P}}^{(L)}=\alpha\left(\mathbf{I}-(1-\alpha)\mathbf{D}^{-1}\mathbf{A^{\prime}}\right)^{-1}.$ (5)
Here, $\mathcal{V}_{L}$ denotes the set of labeled nodes, and $\tilde{P}^{(L)}_{v,u}$ represents the influence of node $v$ on node $u$ as captured by $\mathbf{\tilde{P}}^{(L)}$. To accurately capture the influence of labeled nodes, the inclusion of self-loops in $\mathbf{A^{\prime}}$ ensures that each labeled node’s own label contributes to its $\Lambda^{l}_{u}$. Compared to the intuitive approach of directly selecting labeled nodes, the SALC metric $\Lambda^{\text{SALC}}_{u}$ considers both structural representativeness of the nodes and diffusion of label signals, thus avoiding biases caused by isolated labeled nodes. It is also capable of selecting unlabeled nodes that still possess rich label signals and structural advantages. This improves knowledge quality, enriches the repository, and mitigates the label signal disruption problem. After computing the SALC scores, we rank the nodes based on their $\Lambda^{\text{SALC}}_{u}$ values and select the top $K$ nodes, where the default value of $K$ is $1/3$ of the total number of nodes. Subsequently, for each class $c$, the local prototype $\mathbf{H}_{c}^{i}$ at each client is computed as the mean feature vector of the selected nodes belonging to class $c$:
@@
where $\mathcal{V}_{c}^{i}$ represents the set of selected nodes categorized as class $c$ on client $i$, and $\mathbf{h}_{u}^{i}$ is the feature vector of node $u$ on client $i$. Once the local prototypes are computed, clients upload their prototypes to the server along with the node count. For each class $c$, the server aggregates the prototypes from $\alpha$ percent of the clients by weighting each local prototype according to its sample size. Four global anchor prototypes are constructed for each class. Each global prototype $\mathbf{H}_{c}^{g,k}$ is computed as:
@@
where $\mathbf{H}_{c}^{g,k}$ represents the $k$-th global prototype for class $c$, $\mathcal{N}_{c}^{k}$ is the set of clients randomly selected to contribute to the $k$-th global prototype for class $c$, and $C$ is the number of classes. The global repository $\mathcal{H}$ contains all the global prototypes and will be broadcast back to clients. After the global knowledge repository is constructed, it is distributed to the local clients along with the model parameters. The global features $\mathbf{h}_{u}^{g}$ used in the following loss formulation refer to the frozen inference features extracted locally using the distributed global model. To regulate local training, we propose a federated knowledge distillation loss function aimed at harmonizing the semantic feature localization of the local GNN with its global counterpart, namely by aligning their similarity distributions for the representative prototypes stored in the global repository:
@@
where $\text{KL}(\cdot,\cdot)$ represents the Kullback-Leibler divergence, and $\sigma(\cdot)$ is the softmax function applied to similarity scores computed by the function $\varphi(\mathbf{h},\mathcal{H})$, which returns a vector of cosine similarities between the feature and all prototypes.
@@ 4.3 Frequency-aware Graph Modeling Alignment
To emphasize the GNN spectral adjacency awareness and more accurately capture similarities between nodes, we leverage the feature matrix $\mathbf{h}$ to compute node similarity matrices. In the construction of the following similarity matrix, $\mathbf{h}$ denotes operations applied to both $\mathbf{h}^{i}$ and $\mathbf{h}^{g}$. Specifically, for each node $u$, we identify its $k_{\text{sim}}$ most similar neighbors based on the cosine similarity of their feature vectors $\mathbf{h}_{u}$. We then construct a sparse self-similarity matrix $\mathbf{S^{\prime}}$ as:
@@
Subsequently, we calculate the graph laplacian matrix $\mathbf{L^{\prime}}$ based on this sparse similarity matrix $\mathbf{S^{\prime}}$ as:
@@
where $\mathbf{D^{\prime}}$ is the diagonal degree matrix of $\mathbf{S^{\prime}}$. Moreover, we perform eigendecomposition on the laplacians $\mathbf{L^{\prime}}^{i}$ and $\mathbf{L^{\prime}}^{g}$. For $\mathbf{L^{\prime}}^{i}$, let $\{\mathbf{u}^{\text{low},i}_{m}\}_{m=1}^{k_{\text{eig}}}$ be the eigenvectors corresponding to the smallest eigenvalues, which represents low-frequency. While $\{\mathbf{u}^{\text{high},i}_{m}\}_{m=1}^{k_{\text{eig}}}$ denotes the largest eigenvalues, which represents high-frequency. Similarly, for $\mathbf{L^{\prime}}^{g}$, we obtain $\{\mathbf{u}^{\text{low},g}_{m}\}_{m=1}^{k_{\text{eig}}}$ and $\{\mathbf{u}^{\text{high},g}_{m}\}_{m=1}^{k_{\text{eig}}}$. The feature matrix $\mathbf{h}$ is then projected onto each of these eigenvectors. For instance:
@@
Applying these projections for each $m\in\{1,\dots,k_{\text{eig}}\}$, we obtain several sets of projected feature matrices used in the loss computation. Specifically, local features $\mathbf{h}^{i}$ are projected onto low/high-frequency eigenvectors of the local graph, yielding $\mathbf{Z}^{i,\text{low}}_{m}$ and $\mathbf{Z}^{i,\text{high}}_{m}$, respectively. Similarly, frozen global inference features $\mathbf{h}^{g}$ are projected onto corresponding eigenvectors from $\mathbf{L^{\prime}}^{g}$, yielding $\mathbf{Z}^{g,\text{low}}_{m}$ and $\mathbf{Z}^{g,\text{high}}_{m}$. Conequently, loss $\mathcal{L}_{\text{{FGMA}}}$ is then defined as the sum of MSE over all eigenvector-projected pairs:
@@
This loss addresses spectral heterogeneity by aligning client and global signal characteristics in both spectral domains. Combining strategies Node Label Information Reinforcementand Frequency-aware Graph Modeling Alignment, our framework $S^{2}$FGL reinforces semantic knowledge during local modeling and mitigates spectral client drift. Ultimately, the loss for local training is:
@@
where $\mathcal{L}_{\text{CE}}$ denotes the standard cross-entropy loss for node classification, while $\lambda_{1}$ and $\lambda_{2}$ are balancing hyperparameters for the proposed methods NLIR and FGMA.
@@ 5.1 Experimental Setup
Table 1: Performance Comparison with the state-of-the-art methods on homophilic and heterophilic graph datasets. We report the node classification accuracies with the performance improvement over FedAvg. The best results are highlighted in bold.
@@
FedAvg [ASTAT17] $81.9\pm 0.7$ $74.3\pm 0.4$ $87.3\pm 0.3$ $2.8\pm 2.2 77$$7.6\pm 2.7$ $79.6\pm 0.1$
FedProx [arXiv18] $82.1\pm 0.5$ ↑0.2 $74.4\pm 0.3$ ↑0.1 $87.9\pm 0.4$ ↑0.6 $73.5\pm 3.7$ ↑0.7 $77.3\pm 3.4$ ↓0.3 $79.7\pm 0.1$ ↑0.1
FedNova [NeurIPS20] $81.6\pm 1.2$ ↓0.3 $74.4\pm 0.4$ ↑0.1 $88.2\pm 0.5$ ↑0.9 $73.0\pm 4.4$ ↑0.2 $77.4\pm 4.2$ ↓0.2 $79.9\pm 0.4$ ↑0.3
FedFa [ICLR23] $82.7\pm 0.5$ ↑0.8 $74.9\pm 0.6$ ↑0.6 $87.8\pm 0.5$ ↑0.5 $73.9\pm 3.6$ ↑1.1 $78.1\pm 4.6$ ↑0.5 $80.1\pm 0.3$ ↑0.5
FedSage+ [NeurIPS19] $82.3\pm 0.7$ ↑0.4 $75.2\pm 0.3$ ↑0.9 $88.2\pm 0.7$ ↑0.9 $73.7\pm 4.0$ ↑0.9 $\mathbf{79.0\pm 3.3}$ ↑1.4 $79.9\pm 0.2$ ↑0.3
FedStar [AAAI23] $82.6\pm 0.5$ ↑0.7 $74.5\pm 0.3$ ↑0.2 $88.1\pm 0.6$ ↑0.8 $74.3\pm 2.7$ ↑1.5 $78.3\pm 4.7$ ↑0.7 $79.8\pm 0.1$ ↑0.2
FedPub [ICML23] $82.3\pm 0.8$ ↑0.4 $74.8\pm 0.7$ ↑0.5 $88.0\pm 0.4$ ↑0.7 $73.4\pm 3.5$ ↑0.6 $77.8\pm 3.1$ ↑0.2 $79.9\pm 0.2$ ↑0.3
FGSSL [IJCAI23] $82.6\pm 0.4$ ↑0.7 $74.9\pm 0.2$ ↑0.6 $87.6\pm 0.7$ ↑0.3 $73.6\pm 4.6$ ↑0.8 $77.8\pm 3.8$ ↑0.2 $79.9\pm 0.2$ ↑0.3
FedGTA [VLDB24] $82.4\pm 0.8$ ↑0.5 $75.1\pm 0.5$ ↑0.8 $87.7\pm 0.9$ ↑0.4 $72.6\pm 4.2$ ↓0.2 $77.8\pm 4.1$ ↑0.2 $80.2\pm 0.3$ ↑0.6
FGGP [AAAI24] $82.5\pm 0.4$ ↑0.6 $74.7\pm 0.5$ ↑0.4 $87.5\pm 0.4$ ↑0.2 $73.6\pm 2.8$ ↑0.8 $78.2\pm 3.4$ ↑0.6 $80.4\pm 0.3$ ↑0.8
$S^{2}$FGL (ours) $\mathbf{83.4\pm 0.2}$ ↑1.5 $\mathbf{76.0\pm 0.3}$ ↑1.7 $\mathbf{88.6\pm 0.2}$ ↑1.3 $\mathbf{74.8\pm 2.3}$ ↑2.0 $\mathbf{79.0\pm 1.0}$ ↑1.4 $\mathbf{80.5\pm 0.1}$ ↑0.9
Datasets. We conducted experiments on various datasets to validate the superiority of our framework $S^{2}$FGL. The homophilic graph datasets include Cora, Citeseer, and Pubmed, while the heterophilic graph datasets comprise Texas, Wisconsin, and Minesweeper. The following provides a description of each dataset. Cora (McCallum et al., 2000) dataset consists of 2708 scientific publications classified into one of seven classes. There are 5429 edges in the network of citations. 1433 distinct words make up the dictionary. Citeseer (Giles et al., 1998) dataset consists of 3312 scientific publications classified into one of six classes and 4732 edges. The dictionary contains 3703 unique words. Pubmed (Sen et al., 2008) dataset consists of 19717 scientific papers on diabetes that have been categorized into one of three categories in the PubMed database. The citation network has 44338 edges. Texas and Wisconsin datasets are subsets of the WebKB dataset (Craven et al., 1998). The WebKB dataset was introduced in 1998, comprising web pages from the computer science departments of various universities, including the University of Texas and the University of Wisconsin. The dataset is commonly used for tasks such as webpage classification and link prediction, serving as a benchmark for evaluating machine learning models in graph-based learning scenarios. Minesweeper (Baranovskiy et al., 2023) dataset is a synthetic graph dataset inspired by the Minesweeper game. In this dataset, the graph is structured as a regular 100x100 grid, where each node represents a cell connected to its neighboring nodes, except for edge nodes, which have fewer neighbors. The primary task is to predict which nodes contain mines. This dataset is commonly used to evaluate the performance of GNNs under heterophily.
Evaluation Metric. Following mainstream FGL research experimantal practices, we utilize the accuracy of the node classification task as the evaluation metric.
@@
Baselines. We compare $S^{2}$FGL with several state-of-the-art approaches, including traditional federated learning methods such as FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), FedNova (Wang et al., 2020), and FedFa (Zhou & Konukoglu, 2023); federated graph learning approaches including FGSSL (Huang et al., 2023a) and FGGP (Wan et al., 2024a); as well as personalized federated graph learning methods such as FedSage+ (Zhang et al., 2021b), FedStar (Tan et al., 2023), FedPub (Baek et al., 2023), and FedGTA (Li et al., 2024). This comprehensive set of baseline methods spans various FL and FGL paradigms, allowing us to evaluate the performance and advantages of our proposed $S^{2}$FGL across diverse scenarios.
@@
Implement Details. Following prevalent methodologies in FGL research, we employ the Louvain community detection algorithm to partition the graph into subgraphs assigned to different clients. For each dataset, we divide the nodes into training, validation, and testing sets with ratios of 60%, 20%, and 20%, respectively. Additionally, we simulate various collaborative scenarios by configuring the number of clients to 10 for Cora, Citeseer, Pubmed, and Minesweeper datasets, and 3 for the Texas and Wisconsin datasets. The primary evaluation metric is the node classification accuracy on the clients’ test sets. We conduct each experiment five times and report the average accuracy from the last five communication epochs as the final performance. We conduct experiments with the ACM-GCN (Luan et al., 2022), which achieves a strong ability on both homophilic and heterophilic graph datasets.